# No Silver Bullets: Why Understanding Software Cycle Time is Messy, Not Magic

### v3

@@ 6.2 General Discussion
Another strength that is relevant to both researchers and practitioners is the inclusion, simultaneously, of multiple factors in the same models. Splitting the data into groups using some factor like coding days and taking the average cycle time for each group, as a basis for making inferences about differences between groups, is not sufficient to account for the possibility that other factors systematically differ between those groups. Nevertheless, this is an approach which we have frequently seen in the industry. Modeling these factors simultaneously as continuous variables allows us to get closer to the idea of “X has an effect on Y, all else being held equal.” This is necessary for starting to figure out the effect of a particular factor in isolation. Of course, if one has access to even more sophisticated observational-causal-inference methods (e.g., natural experiments, or matching) or even interventional approaches, these would be even better. If this sounds a little bit difficult, it’s because, like many things in software development, it is. There’s no free lunch; and there are no solutions, only tradeoffs. If one wants easy quantitative answers, one must be willing to trade-off accuracy (and we almost never can say how much accuracy is lost until we do it the hard way too).
@@ 7 Acknowledgements
8 Contributions
Contribution roles listed are defined in NISO CRediT Working Group (2022).

### v4

@@ 6.2 General Discussion
Another strength that is relevant to both researchers and practitioners is the inclusion, simultaneously, of multiple factors in the same models. Splitting the data into groups using some factor like coding days and taking the average cycle time for each group, as a basis for making inferences about differences between groups, is not sufficient to account for the possibility that other factors systematically differ between those groups. Nevertheless, this is an approach which we have frequently seen in the industry. While not necessarily uncommon in this literature, it is worth mentioning that modeling these factors simultaneously as continuous variables allows us to get closer to the idea of “X has an effect on Y, all else being held equal.” This is necessary for starting to figure out the effect of a particular factor in isolation. Of course, if one has access to even more sophisticated observational-causal-inference methods (e.g., natural experiments, or matching) or even interventional approaches, these would be even better. If this sounds a little bit difficult, it’s because, like many things in software development, it is. There’s no free lunch; and there are no solutions, only tradeoffs. If one wants easy quantitative answers, one must be willing to trade-off accuracy (and we almost never can say how much accuracy is lost until we do it the hard way too).
@@ 7 Acknowledgements
8 Declarations
8.1 Ethical approval
This research used aggregated, anonymized GitHub activity data routinely collected through our company’s normal operations and permitted by our Terms of Service. No personal information was gathered specifically for this study, and strict protocols were followed to prevent re-identification of individuals or organizations. Because the dataset was pre-existing, fully anonymized, and did not involve direct interaction with human subjects, the research is exempt from IRB review under 45 CFR §46.104(d)(4)(ii). All data was stored on secure systems with limited access, ensuring both data integrity and confidentiality.
8.2 Author Contributions
Contribution roles listed are defined in Group (2022).
@@ 8.2 Author Contributions
8.3 Data Availability Statement
Data are considered proprietary and are not available to be shared. Code for these analyses is available as analyses.qmd, here: https://github.com/jflournoy/no-silver-bullets.
8.4 Conflict of Interest
The authors are employed by Pluralsight, which is the source of the data used in this study. However, the authors’ roles do not involve product sales or marketing, and this research does not evaluate product performance. The company approved the use of anonymized data for this research but did not influence the analysis methods, findings, or conclusions presented in this paper.
8.5 Other declarations
  • Funding: Not applicable.
  • Informed consent: Not applicable.
  • Clinical trial number: not applicable.